# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
$\geq\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\Phi_{\ell\beta_{\gamma,n},j_{0}}\bigg{(}(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{)}$
$\sim\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\operatorname{\exp}\bigg{(}\ell\beta_{\gamma,n}\bigg{|}(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{|}^{\frac{n}{n-2}}\bigg{)}$
$=\frac{\omega_{n-1}}{(n-\gamma)}~{}\operatorname{\exp}\bigg{[}\bigg{(}1-\frac{\gamma}{n}\bigg{)}\ln k\bigg{(}\frac{1+\varepsilon}{(1+\zeta_{k})^{\frac{n}{n-2}}}-1\bigg{)}\bigg{]}\to+\infty\quad\text{as}\quad k\to\infty,$
from which the proof of Theorem 1.4 is well established. ∎
4. Proof of Theorem 1.5 and Theorem 1.6
Basically, the purpose of this section is to prove Theorem 1.5 and Theorem 1.6. More precisely, we shall demonstrate the asymptotic behavior for the supremum $\text{ATSC}(\ell,\gamma)$ for the subcritical sharp singular Adams’ type inequality. In fact, we also mention here that the proof of Theorem 1.5 doesn’t require the validity of $\text{ATC}_{\frac{n}{2},\frac{n}{2}}(\gamma)<+\infty$. Additionally, we also provide the relationship between the suprema for the critical and subcritical sharp singular Adams’ type inequalities. Now we establish the following crucial lemmas, which are useful in the proof of Theorem 1.5 and Theorem 1.6.
@@
Let $u\in E$ be such that $\|\Delta u\|_{\frac{n}{2}}\leq 1$. Define
@@
$=\cfrac{1}{\|\Delta u\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x,$
from which we conclude the proof. ∎
Lemma 4.2.
Proof.
Let $u\in E$ be such that $\|\Delta u\|_{\frac{n}{2}}\leq 1$ and $\|\Delta u\|_{p}=1$. Define
$w(x)=\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{n}}u(\lambda x)\quad\text{with}\quad\lambda=\Bigg{(}\cfrac{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}\Bigg{)}^{\frac{p}{b(n-2p)}}.$
By direct calculations, we have
$\|\Delta w\|^{a}_{\frac{n}{2}}=\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\big{(}\frac{n-2}{n}\big{)}a}\|\Delta u\|^{a}_{\frac{n}{2}}\leq\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\big{(}\frac{n-2}{n}\big{)}a}$
and
$\|\Delta w\|^{b}_{p}$ $=\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{(\frac{n-2}{n})b}\lambda^{-b\big{(}\frac{n-2p}{p}\big{)}}\|\Delta u\|^{b}_{p}=1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{(\frac{n-2}{n})a}.$
It follows that $\|\Delta w\|^{a}_{\frac{n}{2}}+\|\Delta w\|^{b}_{p}\leq 1$. Consequently, we deduce from the definition of $\text{ATC}_{a,b}(\gamma)$ that
$\leq\Bigg{(}\cfrac{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATC}_{a,b}(\gamma),$
from which the proof of Lemma 4.2 is well established. ∎
Note that Lemma 4.1 allows us to assume $\|\Delta u\|_{p}=1$ in the subcritical sharp singular Adams’ type inequality. Now we can prove Theorem 1.5.
@@
On exploiting the denseness of $C_{0}^{\infty}(\mathbb{R}^{n})$ in $E$ and Lemma 4.1, without loss of generality, one can assume $u\in C_{0}^{\infty}(\mathbb{R}^{n})\setminus\{0\}$, $\|\Delta u\|_{\frac{n}{2}}\leq 1$ and $\|\Delta u\|_{p}=1$. Further, since $|u|=\sqrt{u^{2}}$, we can easily deduce that $|\Delta|u||=|\Delta u|$. Let us define
$S:=\Bigg{\{}x\in\mathbb{R}^{n}\colon~{}|u(x)|>\bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n}}\bigg{\}}.$
Now using again the denseness of $C_{0}^{\infty}(\mathbb{R}^{n})$ in $D^{2,p}(\mathbb{R}^{n})$ and the embedding $D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$, one has that $S$ is a bounded domain. Indeed, we have
$|S|=\int_{S}1\,\mathrm{d}x$ $\leq\cfrac{1}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}}}~{}\|u\|^{p^{\ast}}_{p^{\ast}}\leq\cfrac{C(n,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}}},$
where $C(n,p)>0$ is a constant depending only on $n$ and $p$. Now join us in writing
@@
To evaluate $I_{2}$, we first notice that $\mathbb{R}^{n}\setminus S\subset\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1\}$. By using the embedding $D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$ and the fact that $\|\Delta u\|_{p}=1$, we have
$I_{2}$$\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\big{(}\ell(1-\frac{\gamma}{n})\big{)}^{j}}{j!}|u|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\big{(}\ell(1-\frac{\gamma}{n})\big{)}^{j}}{j!}|u|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1,~{}|x|<1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|u(x)|\leq 1,~{}|x|\geq 1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|\geq 1\}}|u|^{p^{\ast}}\,\mathrm{d}x\leq C(n,\gamma,p),$ (4.1)
where $C(n,\gamma,p)>0$ is a constant depending only on $n$, $\gamma$ and $p$. Finally, to evaluate $I_{1}$, we define
@@
It follows that $w\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and $\|\Delta w\|_{\frac{n}{2}}\leq 1$. By employing (3.6) with $\eta=\frac{\beta(n,2)}{\ell}-1$, we have
$|u|^{\frac{n}{n-2}}$ $\leq\Bigg{(}|w|+\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n}}\Bigg{)}^{\frac{n}{n-2}}\leq(1+\eta)|w|^{\frac{n}{n-2}}+C_{\eta}\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n-2}}$
$=\frac{\beta(n,2)}{\ell}|w|^{\frac{n}{n-2}}+\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{-\frac{2}{n-2}}\Bigg{(}1-\Big{(}\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{n-2}{2}}\Bigg{)}^{\frac{2}{n-2}}$
$=\frac{\beta(n,2)}{\ell}|w|^{\frac{n}{n-2}}+1.$
Gathering all the above information, we obtain from the singular Adams’ inequality [32, Theorem 1.2] with homogeneous Navier boundary condition that
$I_{1}$ $\leq\int_{S}\frac{\operatorname{\exp}(\ell(1-\frac{\gamma}{n})|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}\int_{S}\frac{\operatorname{\exp}(\beta_{\gamma,n}|w|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg{(}\ell\bigg{(}1-\frac{\gamma}{n}\bigg{)}\bigg{)}C(\gamma,n)|S|^{1-\frac{\gamma}{n}}\leq\cfrac{C(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}},$ (4.2)
where $C(n,\gamma,p)>0$ is a constant depending only on $n$, $\gamma$ and $p$. Hence we obtain from (4) and (4) that
$\text{ATSC}(\ell,\gamma)\leq\cfrac{C(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}.$
@@
$v(s)=\Psi\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)},~{}\forall~{}s>0\quad\text{and}\quad r\in(0,1).$ (4.3)
@@
where $\alpha(n,2)=\omega_{n-1}(n-2)$ and the big $O$ term involves terms that are constant mutiple of $\Psi^{\prime\prime}\Big{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\Big{)}\big{(}\log\frac{1}{r}\big{)}^{-2}$. In addition, by direct computations, we also have
@@
where we used the following estimates
$\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{r^{1-\varepsilon}}^{r^{\varepsilon}}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}\leq 1;$
$\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{a}^{b}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}=0,~{}\forall~{}(a,b)\in\{(0,r),(1,+\infty)\}$
and
$\Big{(}\log\frac{1}{r}\Big{)}^{-1}\int_{a}^{b}\Bigg{|}(-1)\Psi^{\prime}\bigg{(}\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg{)}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Bigg{|}^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}\leq$ $~{}\varepsilon\Big{(}\|f^{\prime}\|_{\infty}+O\Big{(}\Big{(}\log\frac{1}{r}\Big{)}^{-1}\Big{)}\Big{)}^{\frac{n}{2}},$
$~{}\forall~{}(a,b)\in\{(r,r^{1-\varepsilon}),(r^{\varepsilon},1)\}.$
@@
where the last inequality is obtained by using the elementary inequality $\frac{x-1}{x}\leq\log x\leq x-1,~{}\forall~{}x>0$. In virtue of the above information, one has
$\text{ATSC}((\beta(n,2),\gamma)$ $\geq\cfrac{1}{\Big{\|}\Delta\Big{(}\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big{)}\Big{\|}_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<r\}}\cfrac{\Phi_{\beta_{\gamma,n},j_{0}}\Big{(}\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big{)}}{|x|^{\gamma}}\,\mathrm{d}x$
@@
This shows the validity of the sharpness of $\beta(n,2)$ and hence the claim is well established. Next, we consider the same sequence $\{\xi_{k}\}_{k}\subset E$ that is used in the proof of Theorem 1.3. Observe that the following estimates hold
@@
where $C_{1}$ and $C_{2}$ are two positive constants.
Define
$w_{k}=\frac{\xi_{k}}{\|\Delta\xi_{k}\|_{\frac{n}{2}}}.$
@@
$\text{ATSC}(\ell,\gamma)$ $\geq\cfrac{1}{\|\Delta w_{k}\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<k^{-\frac{1}{n}}\}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(w_{k})}{|x|^{\gamma}}\,\mathrm{d}x$
@@
$\text{ATSC}(\ell,\gamma)\gtrsim(\ln k)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}~{}\operatorname{\exp}\Big{[}\Big{(}1-\frac{\gamma}{n}\Big{)}\ln k\Big{(}\frac{\ell}{\beta(n,2)}-1\Big{)}\Big{]}.$
@@
$\text{ATSC}(\ell,\gamma)\gtrsim\cfrac{\operatorname{\exp}\Big{(}-\Big{(}1-\frac{\gamma}{n}\Big{)}p\Big{)}}{\Big{(}1-\frac{\ell}{\beta(n,2)}\Big{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}\thickapprox\cfrac{\frac{n-2}{2}~{}\operatorname{\exp}\Big{(}-\Big{(}1-\frac{\gamma}{n}\Big{)}p\Big{)}}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}},$
where we used
@@
$\text{ATSC}(\ell,\gamma)\geq\cfrac{c(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}},$

### v3

@@ 2. Preliminary results
$\geq\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~\Phi_{\ell\beta_{\gamma,n},j_{0}}\Bigg((1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\bigg(\frac{\ln k}{\beta(n,2)}\bigg)^{1-\frac{2}{n}}\Bigg)$
$\sim\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~\operatorname{\exp}\Bigg(\ell\beta_{\gamma,n}\Bigg|(1+\zeta_{k})^{-1}(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\bigg(\frac{\ln k}{\beta(n,2)}\bigg)^{1-\frac{2}{n}}\Bigg|^{\frac{n}{n-2}}\Bigg)$
$=\frac{\omega_{n-1}}{(n-\gamma)}~\operatorname{\exp}\Bigg[\bigg(1-\frac{\gamma}{n}\bigg)\ln k\Bigg(\frac{1+\varepsilon}{(1+\zeta_{k})^{\frac{n}{n-2}}}-1\Bigg)\Bigg]\to+\infty\quad\text{as}\quad k\to\infty.$
This finishes the proof of Theorem 1.2. ∎
5. Asymptotic behavior of subcritical sharp singular Adams’ type Inequality: Proof of Theorem 1.3
Basically, the purpose of this section is to prove Theorem 1.3. More precisely, we shall demonstrate the asymptotic behavior of the supremum $\text{ATSC}(\ell,\gamma)$ for the subcritical sharp singular Adams’ type inequality. To prove this, we establish the following crucial result.
@@
Let $u\in E$ be such that $\|\Delta u\|_{\frac{n}{2}}\leq 1$. Define the function $w$ as follows
@@
$=\|\Delta u\|_{p}^{-p^{\ast}(1-\frac{\gamma}{n})}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x.$
Hence, we infer from the above equality that the proof of Lemma 5.1 is completed. ∎
Due to Lemma 5.1, one can assume $\|\Delta u\|_{p}=1$ in the subcritical sharp singular Adams’ type inequality. Now, we finish this section by proving Theorem 1.3 as follows.
@@
By exploiting the denseness of $C_{0}^{\infty}(\mathbb{R}^{n})$ in $E$ and Lemma 5.1, without loss of generality, one can assume $u\in C_{0}^{\infty}(\mathbb{R}^{n})\setminus\{0\}$, $\|\Delta u\|_{\frac{n}{2}}\leq 1$ and $\|\Delta u\|_{p}=1$. Further, since $|u|=\sqrt{u^{2}}$, we can easily deduce that $|\Delta|u||=|\Delta u|$. Define the set $S$ as follows
$S:=\Bigg\{x\in\mathbb{R}^{n}\colon~|u(x)|>\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{\frac{2}{n}}\Bigg\}.$
Now, using again the denseness of $C_{0}^{\infty}(\mathbb{R}^{n})$ in $D^{2,p}(\mathbb{R}^{n})$ and the embedding $D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$, one sees that $S$ is a bounded domain. Indeed, we have
$|S|=\int_{S}1\,\mathrm{d}x$ $\leq\|u\|^{p^{\ast}}_{p^{\ast}}\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{-\frac{2p^{\ast}}{n}}\leq C(p,n)\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{-\frac{2p^{\ast}}{n}},$
where $C(p,n)>0$ is a constant depending only on $p$ and $n$. Next, join us in writing
@@
To evaluate $I_{2}$, we first notice that $\mathbb{R}^{n}\setminus S\subset\{x\in\mathbb{R}^{n}\colon~|u(x)|\leq 1\}$. By using the embedding $D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$ and the fact that $\|\Delta u\|_{p}=1$, we get
(5.1) $I_{2}$ $\leq\int_{\{x\in\mathbb{R}^{n}\colon~|u(x)|\leq 1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\big(\ell(1-\frac{\gamma}{n})\big)^{j}}{j!}|u|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{\{x\in\mathbb{R}^{n}\colon~|u(x)|\leq 1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\big(\ell(1-\frac{\gamma}{n})\big)^{j}}{j!}|u|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg(\ell\bigg(1-\frac{\gamma}{n}\bigg)\bigg)\Bigg[\int_{\{x\in\mathbb{R}^{n}\colon~|u(x)|\leq 1,~|x|<1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in\mathbb{R}^{n}\colon~|u(x)|\leq 1,~|x|\geq 1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x\Bigg]$
$\leq\operatorname{\exp}\bigg(\ell\bigg(1-\frac{\gamma}{n}\bigg)\bigg)\Bigg[\int_{\{x\in\mathbb{R}^{n}\colon~|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in\mathbb{R}^{n}\colon~|x|\geq 1\}}|u|^{p^{\ast}}\,\mathrm{d}x\Bigg]$
$\leq C(\gamma,p,n),$
where $C(\gamma,p,n)>0$ is a constant depending only on $\gamma$, $p$ and $n$. Finally, to evaluate $I_{1}$, we define
@@
It follows that $w\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and $\|\Delta w\|_{\frac{n}{2}}\leq 1$. By employing (3.6) with $\eta=\frac{\beta(n,2)}{\ell}-1$, we obtain
$|u|^{\frac{n}{n-2}}$ $\leq\Bigg(|w|+\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{\frac{2}{n}}\Bigg)^{\frac{n}{n-2}}\leq(1+\eta)|w|^{\frac{n}{n-2}}+C_{\eta}\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{\frac{2}{n-2}}$
$=\frac{\beta(n,2)}{\ell}|w|^{\frac{n}{n-2}}+\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{-\frac{2}{n-2}}\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{\frac{2}{n-2}}$
$=\frac{\beta(n,2)}{\ell}|w|^{\frac{n}{n-2}}+1\quad\text{in}\quad S.$
By gathering all the above information, we obtain from the singular Adams’ inequality [32, Theorem 1.2] with homogeneous Navier boundary conditions that
(5.2) $I_{1}$ $\leq\int_{S}\frac{\operatorname{\exp}\big(\ell(1-\frac{\gamma}{n})|u|^{\frac{n}{n-2}}\big)}{|x|^{\gamma}}\,\mathrm{d}x\leq\operatorname{\exp}\bigg(\ell\bigg(1-\frac{\gamma}{n}\bigg)\bigg)\int_{S}\frac{\operatorname{\exp}\big(\beta_{\gamma,n}|w|^{\frac{n}{n-2}}\big)}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}\bigg(\ell\bigg(1-\frac{\gamma}{n}\bigg)\bigg)C(\gamma,n)|S|^{1-\frac{\gamma}{n}}\leq C(\gamma,p,n)\bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\bigg)^{-\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)},$
where $C(\gamma,p,n)>0$ is a constant depending only on $\gamma$, $p$ and $n$. Hence, we obtain from (5.1) and (5.2) that
$\text{ATSC}(\ell,\gamma)\leq C(\gamma,p,n)\bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\bigg)^{-\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}.$
@@
(5.3) $v(s)=\Psi\bigg(\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg),~\forall~s>0\quad\text{and}\quad r\in(0,1).$
@@
where $\alpha(n,2)=\omega_{n-1}(n-2)$ and the big $O$ term involves terms that are constant multiple of $\Psi^{\prime\prime}\Big(\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\Big)\big(\log\frac{1}{r}\big)^{-2}$. In addition, by direct computations, we also have
@@
where we have used the following estimates
  • (a)
    There holds
    $\Big(\log\frac{1}{r}\Big)^{-1}\int_{r^{1-\varepsilon}}^{r^{\varepsilon}}\Bigg|(-1)\Psi^{\prime}\bigg(\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg)+O\Big(\Big(\log\frac{1}{r}\Big)^{-1}\Big)\Bigg|^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}\leq 1,$
  • (b)
    For all $(a,b)\in\{(0,r),(1,+\infty)\}$, there holds
    $\Big(\log\frac{1}{r}\Big)^{-1}\int_{a}^{b}\Bigg|(-1)\Psi^{\prime}\bigg(\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg)+O\Big(\Big(\log\frac{1}{r}\Big)^{-1}\Big)\Bigg|^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}=0,$
  • (c)
    For all $(a,b)\in\{(r,r^{1-\varepsilon}),(r^{\varepsilon},1)\}$, there holds
    $\Big(\log\frac{1}{r}\Big)^{-1}\int_{a}^{b}\Bigg|(-1)\Psi^{\prime}\bigg(\frac{\log\frac{1}{s}}{\log\frac{1}{r}}\bigg)+O\Big(\Big(\log\frac{1}{r}\Big)^{-1}\Big)\Bigg|^{\frac{n}{2}}\,\frac{\mathrm{d}s}{s}\leq$ $~\varepsilon\Big(\|f^{\prime}\|_{\infty}+O\Big(\Big(\log\frac{1}{r}\Big)^{-1}\Big)\Big)^{\frac{n}{2}}.$
@@
where the last inequality is obtained by using the elementary inequality $\frac{x-1}{x}\leq\log x\leq x-1$ for all $x>0$. In virtue of the above information, one has
$\text{ATSC}((\beta(n,2),\gamma)$
$\geq\cfrac{1}{\Big\|\Delta\Big(\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big)\Big\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\{x\in\mathbb{R}^{n}\colon~|x|<r\}}\cfrac{\Phi_{\beta_{\gamma,n},j_{0}}\Big(\frac{\bar{v}}{\|\Delta\bar{v}\|_{\frac{n}{2}}}\Big)}{|x|^{\gamma}}\,\mathrm{d}x$
@@
This shows the validity of the sharpness of $\beta(n,2)$ and thus, the claim is well-established. Next, we consider the same sequence $\{\xi_{k}\}_{k}\subset E$ that is used in the proof of Theorem 1.1. Notice that the following estimates hold
@@
where $C_{1}$ and $C_{2}$ are two positive constants. Now, we define the sequence $\{w_{k}\}_{k}$ as follows
$w_{k}=\frac{\xi_{k}}{\|\Delta\xi_{k}\|_{\frac{n}{2}}},~\forall~k\in\mathbb{N}.$
@@
$\text{ATSC}(\ell,\gamma)$
$\geq\cfrac{1}{\|\Delta w_{k}\|_{p}^{p^{\ast}(1-\frac{\gamma}{n})}}\int_{\{x\in\mathbb{R}^{n}\colon~|x|<k^{-\frac{1}{n}}\}}\frac{\Phi_{\ell(1-\frac{\gamma}{n}),j_{0}}(w_{k})}{|x|^{\gamma}}\,\mathrm{d}x$
@@
$\text{ATSC}(\ell,\gamma)\gtrsim(\ln k)^{\frac{2p^{\ast}}{n}(1-\frac{\gamma}{n})}~\operatorname{\exp}\Bigg[\Big(1-\frac{\gamma}{n}\Big)\ln k\bigg(\frac{\ell}{\beta(n,2)}-1\bigg)\Bigg].$
@@
$\text{ATSC}(\ell,\gamma)\gtrsim\cfrac{\operatorname{\exp}\Big(-\Big(1-\frac{\gamma}{n}\Big)p\Big)}{\Big(1-\frac{\ell}{\beta(n,2)}\Big)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}}\thickapprox\cfrac{\big(\frac{n-2}{2}\big)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}~\operatorname{\exp}\Big(-\Big(1-\frac{\gamma}{n}\Big)p\Big)}{\bigg(1-\big(\frac{\ell}{\beta(n,2)}\big)^{\frac{n-2}{2}}\bigg)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}},$
where we have used
@@
$\text{ATSC}(\ell,\gamma)\geq c(n,\gamma,p)\bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\bigg)^{-\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)},$